Please also help with the last part
/20 points A bus contains a 1,500 kg flywheel (a disk that has a 0.600 m radius) and has a total mass of 10,000 kg. (a) Calculate the angular velocity (in,rad/s) the flywheel must have to contan 10.4 P 027. angular velocity (in, rad/s) the flywheel must have to contan enough energy to take the bus from rest to a speed of 26.0 m/s, assuming 88.0% of the rotational energy , kinetic energy can be transformed into translational x rad/s (b) How h high a hill (in m) can the bus cimb with this stored energy and still heve a speed of 2.40 m/s at the top of the hil? Explicitly show how you folow the steps found in Probiem-Solving Strategy for Rotational Dynamics. (Submit a fie with a maximum size of 1 MB.) Choow Fle No file chosen This answer has not been graded ye Additional Materials Reading
Please also help with the last part
Answer #1

a. KE gained = 0.5mv^2 = 3.38*10^6 J
KE from fly wheel = KE/0.88 = 3.381 *10^6 J = 0.5*I*w^2 = 0.5*m*r^2*w^2
w = 119.271 rad/s
b. From consevation of energy
3.38*10^6 = 0.5*10,000*2.4^2 + 10,000*9.8*h
h = 34.195 m